# FEDONet: Fourier-Embedded DeepONet for Spectrally Accurate Operator Learning

### v3

@@ 2.3 Evaluation and Generalization
    $\varepsilon_{L^{2}}=\frac{\|G_{\theta}(u)-G(u)\|_{2}}{\|G(u)\|_{2}},$
    where the norm is computed over a dense evaluation grid. This metric quantifies the normalized discrepancy between predicted and reference output fields.
  • Spectral Fidelity: We perform Fourier analysis of the predicted and true fields to evaluate preservation of high-frequency content. Specifically, we compare the angle-integrated power spectra $E(k)$ across wavenumbers $k$, assessing the model’s ability to capture multiscale dynamics and resolve fine-scale features.
The ability to evaluate the learned operator at arbitrary coordinates makes FEDONet particularly suitable for scientific computing tasks requiring continuous surrogate modeling. In all benchmark scenarios, FEDONet demonstrates strong generalization, enhanced spectral accuracy, and robust performance across a wide range of PDE systems with varying stiffness, nonlinearity, and complexity.
@@ 3 Results
We constructed a synthetic dataset comprising numerical solutions to the canonical two-dimensional Poisson equation. The governing equation is
@@
where $u(x,y)$ denotes the potential field and $f(x,y)$ represents a prescribed source distribution.
A total of N=10, [^0] N=10{,}000 distinct source fields were synthesized using Gaussian random fields (GRFs) with spectral parameters $\alpha=3$ and $\tau=3$,
@@
on a uniform $128\times 128$ grid. For each realization of $f$, the corresponding solution $u$ was obtained by numerically solving the Poisson equation using a standard five-point finite-difference discretization of the Laplacian operator. The spatial domain $\Omega=[0,1]^{2}$ was thus transformed into a structured algebraic system,
@@
where $\mathbf{u},\mathbf{b}\in\mathbb{R}^{16384}$ denote the vectorized forms of the solution and source fields, respectively. The coefficient matrix $A$ was defined using a conventional five-point stencil with Dirichlet conditions enforced by setting boundary rows to the identity and zeroing the corresponding entries in $\mathbf{b}$. This process yielded 10,000 pairs of input-output fields suitable for training, validation, and testing of neural operator architectures.
@@
Figure 2: Solving a 2D Poisson equation: Comparison of predicted and reference solution fields of DeepONet and FEDONet
@@
Figure 2 shows a representative test case from the Poisson dataset. The top row depicts the input forcing field, exact solution, prediction from the baseline DeepONet, and the corresponding pointwise error, while the bottom row presents results from the proposed Fourier-Embedded DeepONet (FEDONet). Both models were evaluated on $1000$ randomly selected unseen test samples, yielding average relative $L^{2}$ errors of $1.41\%$ for the DeepONet and $1.09\%$ for the FEDONet. FEDONet reproduces the reference field with sharper gradients and more accurate spatial localization of extrema. The residual field of the vanilla DeepONet displays fine-scale artifacts, while FEDONet yields a smoother, spatially coherent error distribution, corresponding to an approximate fivefold improvement in reconstruction accuracy.
Figure 3: Energy spectrum comparison between the ground truth, vanilla and Fourier DeepOnet reconstructions for the 2D Poisson Dataset.
@@
Spectral accuracy was assessed using the angle-integrated energy spectrum $E_{k}$ shown in Fig. 3. Both models reproduce the reference spectrum at large scales ($k\lesssim 50$), but differences arise in the intermediate and dissipative ranges. The baseline DeepONet overestimates energy for $k\gtrsim 90$, resulting in a spurious buildup of high-wavenumber content that mirrors the oscillatory residuals seen in physical space. In contrast, FEDONet closely follows the reference spectrum across the entire range. Taken together, the spatial and spectral analyses demonstrate that the Fourier embedding enhances the trunk network’s ability to capture high-frequency features and improves overall stability. FEDONet achieves lower pointwise errors while eliminating the artificial spectral plateau that characterizes conventional fully connected trunks.
3.2 Burgers’ Equation
The Burgers’ equation is a well-known 1D viscous equation. It has been studied extensively for DeepONets in prior literature [53]. The equation is given by:
where $x$ and $t$ denote the spatio-temporal coordinates and $\nu$ is the kinematic viscosity. We set $\nu=0.01$ along with initial and periodic boundary conditions as:
where the initial condition $s(x)$ is sampled from a Gaussian Random Field, $s(x)\sim\mathcal{N}\left(0,25^{2}(-\Delta+5^{2}I)^{-4}\right)$, satisfying periodicity. The objective is to learn the nonlinear operator $\mathcal{G}$ mapping $s(x)$ to the full spatio-temporal solution $u(x,t)$ using both vanilla DeepONet and the proposed FEDONet.
Both models were evaluated on $250$ unseen test samples, yielding average relative $L^{2}$ errors of $6.43\%$ for the DeepONet and $4.86\%$ for the FEDONet. Figure 4 presents prediction for a test sample, showing the ground truth solution alongside predictions from both models. While DeepONet captures the coarse solution structure, it struggles with steep gradients and exhibits visible smoothing near shock regions. In contrast, FEDONet produces significantly sharper and more accurate reconstructions, faithfully capturing shock fronts and fine-scale dynamics.
Figure 4: Solving a parametric Burgers’ equation: Comparison of predicted spatio-temporal field for a random initial condition. Relative $L^{2}$ error: $4.09\%$ for DeepONet, $1.49\%$ for FEDONet.
The energy spectra of the predicted and ground truth fields, shown in Figure 5, further emphasize this advantage. FEDONet accurately captures the spectral decay across a wide range of wavenumbers, aligning with the physical expectations of dissipative transport phenomena. DeepONet, however, underestimates energy at higher modes, reflecting its inability to resolve fine features and spectral richness. This demonstrates that the Fourier embeddings effectively bridge the model’s spectral bias, enabling it to better represent high-frequency dynamics essential to Burgers-type flows. The enhanced performance of FEDONet in both physical and spectral domains highlights its potential as a more robust and generalizable operator learner for convection-dominated systems.
Figure 5: Energy spectrum and pointwise absolute error across space and time for a representative sample for Burgers’ Equation.
Figure 6 shows the convergence of training losses for different batch sizes. FEDONet consistently converges faster and to lower loss values across all training configurations, indicating improved optimization characteristics and stability during training.
Figure 6: Training loss convergence for the 1D Burgers’ Equation - Comparison of DeepONet and FEDONet trained with varying batch sizes.
3.3 Lorenz-63 System
The Lorenz-63 system is a widely studied chaotic dynamical system that serves as a benchmark for evaluating the long-term stability and predictive accuracy of surrogate models. It is governed by the following system of coupled nonlinear ordinary differential equations:
$\frac{dx}{dt}$ $=\sigma(y-x),$ (15)
$\frac{dy}{dt}$ $=x(\rho-z)-y,$
$\frac{dz}{dt}$ $=xy-\beta z,$
where the canonical parameters are set to $\sigma=10$, $\rho=28$, and $\beta=8/3$. Under these settings, the system exhibits chaotic dynamics on a strange attractor in three-dimensional phase space.
The objective is to approximate the solution operator $\mathcal{G}\colon x_{0}\mapsto\mathbf{x}(t)$, where $\mathbf{x}(t)=(x(t),y(t),z(t))$ denotes the full trajectory over the time interval $t\in[0,3]$, given a scalar initial condition $x_{0}\in[10,15]$, with $y_{0}=z_{0}=12$ held fixed. This configuration reduces the input to a one-dimensional function-to-function mapping, while preserving the rich, sensitive behavior of the output trajectory.
A dataset of 10,00010{,}000 trajectories is generated using the classic Runge-Kutta method (RK45), each evaluated at 1000 uniformly spaced time steps. The input to the DeepONet branch network is the scalar $x_{0}$, and the trunk network is queried at time coordinates $t\in[0,3]$. The goal is to accurately recover the full state trajectory, including lobe transitions and fine-scale oscillations characteristic of chaotic dynamics.
Figure 7: Reconstructed Lorenz-63 attractor for a randomnly sampled initial condition for trained DeepONet and FEDONet
Figure 7 shows the reconstructed attractor for a representative test sample. While DeepONet departs from the ground truth early in the rollout, FEDONet maintains geometric coherence and captures both wings of the attractor. This result demonstrates the stabilizing effect of spectral enrichment on trajectory learning.
Figure 8: Phase-plane projection of the xxzz states.
In Figure 8, we examine the xxzz projection. The DeepONet prediction deviates from the true phase-space trajectory and fails to capture the correct lobe transitions. In contrast, FEDONet successfully recovers the attractor’s phase geometry, indicating more robust learning of long-term dependencies.
Figure 9: Pointwise prediction error in $x(t)$, $y(t)$, and $z(t)$ for a representative sample.
Figure 9 presents the pointwise prediction errors over time. DeepONet accumulates noticeable errors, particularly in the $x(t)$ and $z(t)$ components, which undergo rapid transitions between lobes. These deviations grow sharply after $t>1.5$, illustrating compounding approximation errors. FEDONet exhibits substantially lower and smoother error profiles, with better robustness during nonlinear state transitions.
Figure 10: Cumulative relative error over time.
In Figure 10, we observe that the cumulative error of DeepONet grows exponentially, characteristic of chaotic divergence. In contrast, FEDONet maintains a flatter cumulative error curve, demonstrating enhanced resilience against long-term instability and better preservation of trajectory fidelity.
Figure 11: Spectral energy density of predicted $x(t)$ trajectories compared with the ground truth.
Finally, Figure 11 compares the spectral content of the predicted and reference trajectories. The vanilla DeepONet fails to capture high-frequency modes, resulting in a smoothed trajectory that diverges from the correct dynamics. FEDONet, equipped with randomized Fourier features, preserves the broadband energy spectrum more accurately - an essential factor in recovering the fine-scale structure and temporal richness of chaotic systems.
These results collectively demonstrate that FEDONet offers improved phase alignment, lower pointwise and cumulative error, and superior spectral accuracy in chaotic dynamical systems. The introduction of Fourier embeddings into the trunk network acts as a spectral preconditioner, enabling the model to overcome inductive biases and improve fidelity under sensitivity to initial conditions.
@@ 3.4 Eikonal Equation
We consider the two-dimensional Eikonal equation, which governs the propagation of wavefronts at unit speed and arises in a wide range of applications including geometric optics, level-set methods, and computational geometry. The equation is given by:
@@
where $\mathbf{x}=(x,y)\in\mathbb{R}^{2}$ denotes the spatial coordinates, $\Omega\subset\mathbb{R}^{2}$ is the computational domain, and $\partial\Omega$ is its boundary. The solution $s(\mathbf{x})$ represents the signed distance from a point $\mathbf{x}\in\Omega$ to the boundary $\partial\Omega$, defined as:
@@
with the distance function $d$ defined by:
@@
A NACA 4-digit airfoil is parameterized by its maximum camber $m$, the location of maximum camber $p$, and the maximum thickness $t$. The mean camber line is given by:
@@
The thickness distribution $y_{t}(x)$ is defined as:
@@
The upper and lower surface coordinates of the airfoil are computed as:
@@
Each airfoil is embedded in a $256\times 256$ pixel computational grid and scaled to fit within a central subdomain. A binary mask is created to define the geometry, and the signed distance function is computed using the Euclidean Distance Transform (EDT) as:
@@
where $d_{\text{out}}$ and $d_{\text{in}}$ represent the unsigned distances outside and inside the airfoil boundary, respectively. To ensure numerical stability, the resulting SDF is normalized to the range [1,1][-1,1]:
@@
A total of 1250 distinct airfoil shapes are synthesized by randomly sampling parameters $m\in[0.01,0.09]$, $p\in[0.1,0.7]$, and $t\in[0.1,0.4]$. Of these, 1000 samples are used for training and 250 for testing. This dataset captures a wide range of geometrical variations and curvature profiles, providing a rigorous benchmark for learning PDE solution operators from geometric input. On this benchmark, the baseline DeepONet attains an average relative $L^{2}$ error of 2.15%, while the proposed FEDONet achieves an improved 1.12% error, demonstrating substantially higher reconstruction accuracy.
@@
Figure 12: Solving a parametric Eikonal equation (airfoils): Comparison of predicted signed distance functions (SDFs) for a representative airfoil geometry. Relative $L^{2}$ error: $2.728\times 10^{-2}$ for DeepONet, $0.912\times 10^{-2}$ for FEDONet.
@@
As shown in Figure 12, FEDONet achieves more accurate recovery of the level-set structure. The DeepONet underperforms in regions of high curvature, particularly near the airfoil’s leading and trailing edges, where geometric non-smoothness introduces steep spatial gradients. FEDONet’s use of Fourier embeddings in the trunk network enhances its ability to resolve these features by expanding the function space with oscillatory basis functions aligned with the data’s spectral demands.
@@
These results confirm that FEDONet offers improved resolution of high-gradient and discontinuous regions in PDE fields, even in geometry-conditioned tasks such as the Eikonal equation. This highlights its potential as a general-purpose surrogate for PDEs defined on complex, spatially varying domains.
3.5 Allen–Cahn Equation

### v4

@@ 2.3 Evaluation
    $\varepsilon_{L^{2}}=\frac{\|G_{\theta}(u)-G(u)\|_{2}}{\|G(u)\|_{2}}.$
    where the norm is computed over a dense evaluation grid. This metric quantifies the normalized discrepancy between predicted and reference output fields and serves as the primary measure of reconstruction accuracy.
  • Spectral Fidelity: To evaluate the preservation of multiscale structures, we analyze the Fourier energy spectrum of predicted and reference fields. Specifically, we compare the angle-integrated power spectra $E(k)$ across wavenumbers $k$. This diagnostic provides a stringent test of the model’s ability to recover high-frequency content, which is often underrepresented in operator learning frameworks. Particular emphasis is placed on the decay rate and tail behavior of $E(k)$, which indicate how well fine-scale dynamics are captured.
  • Robustness to Input Noise: We evaluate model performance under varying levels of additive noise in the input functions,
    $u_{\text{test}}^{\sigma}=u_{\text{test}}+\sigma\,\eta,$
    where $\eta$ is Gaussian noise. Here, $\sigma$ is expressed as a fraction of the standard deviation of the input signal, and is reported in percentage terms (e.g., $\sigma=0.1$ corresponds to $10\%$ noise). The degradation of $\varepsilon_{L^{2}}$ with increasing noise levels provides insight into the stability of the learned operator. This analysis is critical for practical scientific applications where measurements are inherently noisy.
  • Sample Efficiency: To assess how efficiently the model learns the operator, we train FEDONet and baseline DeepONet models using varying numbers of training samples. The resulting error scaling with respect to dataset size reveals the data efficiency of the architecture and highlights the impact of Fourier embeddings in low-data regimes.
@@ 3 Results
3.1 Burgers’ Equation
The viscous Burgers equation is a canonical 1D nonlinear PDE. It has been studied extensively for DeepONets in prior literature [56]. We consider the governing equation,
where $x$ and $t$ denote the spatio-temporal coordinates and $\nu$ is the kinematic viscosity. We set $\nu=0.01$ along with initial and periodic boundary conditions as,
where the initial condition $s(x)$ is sampled from a Gaussian Random Field, $s(x)\sim\mathcal{N}\left(0,25^{2}(-\Delta+5^{2}I)^{-4}\right)$, satisfying periodicity. The objective is to learn the nonlinear operator $\mathcal{G}$ mapping $s(x)$ to the full spatio-temporal solution $u(x,t)$ using both vanilla DeepONet and the FEDONet.
Both models are evaluated on $250$ unseen, noise-free test samples, yielding average relative $L^{2}$ errors of $6.43\%$ for DeepONet and $4.86\%$ for FEDONet. Figure 2 shows predictions for a test input corrupted with $10\%$ additive noise. DeepONet captures the coarse solution structure but fails to resolve sharp gradients, exhibiting pronounced smoothing near shock regions. In contrast, FEDONet yields markedly sharper reconstructions, accurately capturing shock fronts and preserving fine-scale dynamics.
Figure 2: Parametric Burgers’ Equation: Comparison of predicted spatio-temporal solution fields for a test input with additive noise. The corresponding relative $L^{2}$ errors are $6.5\%$ for DeepONet and $5.0\%$ for FEDONet.
The energy spectra of the predicted and ground truth fields, shown in Figure 3, further emphasize this advantage. FEDONet accurately captures the spectral decay across a wide range of wavenumbers, aligning with the physical expectations of dissipative transport phenomena. DeepONet, however, underestimates energy at higher modes, reflecting its inability to resolve fine features and spectral richness.
Figure 3: Spectral and Error analysis for Burgers’ equation: Comparison of the energy spectrum and pointwise absolute error fields for a representative noise-free test sample.
Figure 4 illustrates the training loss convergence for DeepONet and FEDONet across a range of batch sizes. Several consistent trends are observed. First, FEDONet exhibits significantly faster initial convergence, rapidly reducing the loss within the early training epochs compared to DeepONet. Second, across all batch sizes, FEDONet consistently attains lower final loss values, indicating a more accurate approximation of the underlying operator. Notably, the performance gap between the two models persists regardless of batch size, demonstrating that the advantage of FEDONet is not merely an artifact of a particular training configuration. As the batch size increases, both models exhibit smoother training curves due to reduced stochasticity; however, DeepONet remains systematically biased toward higher loss values and slower convergence.
Figure 4: Training convergence for Burgers’ equation: Comparison of training loss evolution for DeepONet and FEDONet across different batch sizes (128, 256, 512, 1024).
Figure 5 examines the impact of the number of Fourier features on FEDONet performance. As the number of features increases, the error decreases rapidly initially, indicating improved capacity to capture high-frequency components. Beyond a certain threshold, the performance saturates and exhibits mild fluctuations, suggesting a balance between representational richness and over-parameterization.
Figure 5: Effect of Fourier Feature Resolution: Mean Relative $L^{2}$ error for FEDONet as a function of the number of Fourier features.
@@ 3.1 Burgers’ Equation
The 2D Poisson equation serves as a canonical model for elliptic partial differential equations arising in a wide range of physical systems, including electrostatics, heat conduction, and potential flow. We consider the governing equation,
@@
where $u(x,y)$ denotes the potential field and $f(x,y)$ represents a prescribed source distribution. A total of N=10, [^0] N=10{,}000 distinct source fields were synthesized using Gaussian random fields (GRFs) with smoothness parameter $\alpha=3$ and length scale $\tau=3$.
@@
on a uniform $128\times 128$ grid. For each realization of $f$, the corresponding solution $u$ was obtained by numerically solving the Poisson equation using a standard five-point finite-difference discretization of the Laplacian operator. Discretization over $\Omega=[0,1]^{2}$ yields the structured algebraic system,
@@
where $\mathbf{u},\mathbf{b}\in\mathbb{R}^{16384}$ denote the vectorized forms of the solution and source fields, respectively. The coefficient matrix $A$ was defined using a conventional five-point stencil with Dirichlet conditions enforced by setting boundary rows to the identity and zeroing the corresponding entries in $\mathbf{b}$.
@@
Figure 6: Solving a 2D Poisson equation: Comparison of predicted and reference solution fields of DeepONet and FEDONet for a test input with $10\%$ additive noise. The corresponding relative $L^{2}$ errors are $2.85\%$ for DeepONet and $1.73\%$ for FEDONet.
@@
Figure 6 presents a representative test case from the Poisson dataset. The top row shows the input noisy forcing field, the ground truth solution, the DeepONet prediction, and the corresponding pointwise error, while the bottom row displays the same quantities for FEDONet. Both models were evaluated on $1000$ randomly selected unseen test samples, yielding average relative $L^{2}$ errors of $1.41\%$ for DeepONet and $1.09\%$ for FEDONet under noise-free input conditions. While both models achieve high overall accuracy, FEDONet consistently produces sharper gradients and more precise spatial localization of extrema. In contrast, the error field of DeepONet exhibits fine-scale oscillatory artifacts, indicating limitations in resolving localized structures. FEDONet, however, yields a smoother and more spatially coherent error distribution, reflecting improved reconstruction quality.
Figure 7: Spectral analysis for 2D Poisson equation: Energy spectra of the ground truth solution and the corresponding reconstructions obtained using DeepONet and FEDONet for a representative noise-free test sample.
@@
Spectral accuracy was assessed using the angle-integrated energy spectrum $E_{k}$ shown in Figure 7. Both models reproduce the reference spectrum at large scales ($k\lesssim 50$), but differences arise in the intermediate and dissipative ranges. The baseline DeepONet overestimates energy for $k\gtrsim 90$, resulting in a spurious buildup of high-wavenumber content that mirrors the oscillatory residuals seen in physical space. In contrast, FEDONet closely follows the reference spectrum across the entire range. Taken together, the spatial and spectral analyses demonstrate that the Fourier embedding enhances the trunk network’s ability to capture high-frequency features and improves overall stability. FEDONet achieves lower pointwise errors while eliminating the artificial spectral plateau that characterizes conventional fully connected trunks.
The robustness of the learned operators to noisy inputs is examined in Fig. 8, which shows the variation of mean relative $L^{2}$ error with increasing levels of additive noise. While both models exhibit degradation in performance as noise increases, FEDONet consistently achieves lower error across all noise levels. Moreover, the rate of error growth is significantly smaller for FEDONet, indicating enhanced stability. Notably, even at higher noise levels (up to $20\%$), FEDONet maintains a relatively low error compared to DeepONet, highlighting its ability to preserve the underlying solution structure despite corrupted inputs.
Figure 8: Robustness to input noise for the 2D Poisson equation: Mean relative $L^{2}$ error as a function of input noise level.
@@ 3.3 Eikonal Equation
We consider the 2D Eikonal equation, which governs the propagation of wavefronts at unit speed and arises in a wide range of applications including geometric optics, level-set methods, and computational geometry. We consider the governing equation,
@@
where $\mathbf{x}=(x,y)\in\mathbb{R}^{2}$ denotes the spatial coordinates, $\Omega\subset\mathbb{R}^{2}$ is the computational domain, and $\partial\Omega$ is its boundary. The solution $s(\mathbf{x})$ represents the signed distance from a point $\mathbf{x}\in\Omega$ to the boundary $\partial\Omega$, defined as,
@@
with the distance function $d$ defined by,
@@
A NACA 4-digit airfoil is parameterized by its maximum camber $m$, the location of maximum camber $p$, and the maximum thickness $t$. The mean camber line is given by,
@@
The thickness distribution $y_{t}(x)$ is defined as,
@@
The upper and lower surface coordinates of the airfoil are computed as,
@@
Each airfoil is embedded in a $256\times 256$ pixel computational grid and scaled to fit within a central subdomain. A binary mask is created to define the geometry, and the signed distance function is computed using the Euclidean Distance Transform (EDT) as,
@@
where $d_{\text{out}}$ and $d_{\text{in}}$ represent the unsigned distances outside and inside the airfoil boundary, respectively. To ensure numerical stability, the resulting SDF is normalized to the range [1,1][-1,1],
@@
A total of 1250 distinct airfoil shapes are synthesized by randomly sampling parameters $m\in[0.01,0.09]$, $p\in[0.1,0.7]$, and $t\in[0.1,0.4]$. Of these, 1000 samples are used for training and 250 for testing. This dataset captures a wide range of geometrical variations and curvature profiles, providing a rigorous benchmark for learning PDE solution operators from geometric input. On this benchmark, the baseline DeepONet attains an average relative $L^{2}$ error of 2.15%, while the proposed FEDONet achieves an improved 1.12% error under noise-free input conditions, demonstrating substantially higher reconstruction accuracy.
@@
Figure 9: Solving a parametric Eikonal equation (airfoils): Comparison of predicted signed distance functions (SDFs) for a representative noise-free airfoil geometry. The corresponding relative $L^{2}$ errors are $2.34\%$ for DeepONet and $0.54\%$ for FEDONet.
@@
As shown in Figure 9, FEDONet achieves more accurate recovery of the level-set structure. The DeepONet underperforms in regions of high curvature, particularly near the airfoil’s leading and trailing edges, where geometric non-smoothness introduces steep spatial gradients. FEDONet’s use of Fourier embeddings in the trunk network enhances its ability to resolve these features by expanding the function space with oscillatory basis functions aligned with the data’s spectral demands. This advantage is further reflected in the structure of the error fields, where the DeepONet solution exhibits anisotropic, ray-like artifacts emanating from regions of high curvature, indicative of an uneven representation of spatial frequencies and a directionally biased approximation of the underlying operator. Such behavior is consistent with the spectral bias of standard neural networks, which tend to under-resolve high-frequency components. In contrast, FEDONet produces a more isotropic and spatially coherent error distribution with reduced magnitude, demonstrating improved consistency across scales and a more faithful reconstruction of the solution, particularly in regions dominated by sharp gradients and geometric complexity.
The effect of training dataset size on model performance is shown in Fig. 10. Both models benefit from increased training data; however, FEDONet consistently achieves lower mean relative $L^{2}$ error across all sample sizes. The improvement is particularly significant in the low-data regime, which indicates that the Fourier-embedded architecture is more sample efficient and better suited for learning the nonlinear mapping associated with the Eikonal equation, which involves sharp geometric features and non-smooth solution manifolds.
Figure 10: Sample efficiency for the Eikonal equation: Mean relative $L^{2}$ error as a function of the number of training samples. FEDONet consistently achieves lower error across all regimes, with a pronounced advantage in the low-data setting.
Figure 11 presents qualitative comparisons of model predictions under increasing levels of additive noise in the input mask. As the noise level increases, the DeepONet predictions exhibit pronounced degradation, with large, structured error patterns emerging near regions of high curvature and along characteristic directions. These artifacts manifest as anisotropic streaks and localized error amplification, indicating instability in the learned operator under perturbations. In contrast, FEDONet maintains stable and visually consistent reconstructions across all noise levels. The corresponding error fields remain relatively diffuse and isotropic, with significantly reduced magnitude even at higher noise levels.
Figure 11: Robustness to input noise for the Eikonal equation: Qualitative comparison of predicted fields and corresponding error distributions under increasing noise levels.
@@
These results confirm that FEDONet offers improved resolution of high-gradient and discontinuous regions in PDE fields, even in geometry-conditioned tasks such as the Eikonal equation. This demonstrates that the Fourier embeddings provide enhanced robustness by enabling a more uniform representation of spatial frequencies, thereby preserving the underlying geometric structure of the solution.
3.4 Allen-Cahn Equation